# Jean DELHAYE

### v2

@@ 6. Non-uniqueness of limit profiles
[16, Prop 2.7] Let $L$ be a Gaussian generating functional on $\mathbb{G}=O_{N}^{+}$ or $U_{N}^{+}$ Then, for any $a_{1},...,a_{n}\in\mathcal{O}(\mathbb{G})$ we have
$L(a_{1}...a_{n})=\sum_{1\leq i<j\leq n}L(a_{i}a_{j})\varepsilon(a_{0}...\check{a}_{i}...\check{a}_{j}...a_{n})-(n-2)\sum_{j=1}^{n}L(a_{j})\varepsilon(a_{1}...\check{a}_{j}...a_{n}).$
@@
    $\chi_{w}\longmapsto-{\ell(w)}N^{\ell(w)-1}\alpha-\Big{(}{(p(w)-q(w))^{2}-\ell(w)}\Big{)}N^{\ell(w)-2}\beta,\quad w\hbox{ word on }\ \{\lozenge,\blacklozenge\},$
    for some $\alpha\in\mathbb{C}$ and $\beta\geq 0$ satisfying $\mathrm{Re}(\alpha)\geq\beta/N$.
@@
Let us first note that any Gaussian cocycle $\eta$ defined on a compact quantum group satisfies the relation $\eta\circ S+\eta=0$ (see the beginning of the proof of [16, Thm 3.11]).
Let $(\rho,\eta,L)$ be a Gaussian Schürmann triple on $O_{N}^{+}$. The relation $\eta\circ S=-\eta$ implies that $\eta\equiv 0$ on the central algebra and therefore, we have
$L(ab)=\varepsilon(a)L(b)+L(a)\varepsilon(b),\quad a,b\in\mathcal{O}(O_{N}^{+})_{0}.$
@@
Let $(\rho,\eta,L)$ be a Gaussian Schürmann triple on $U_{N}^{+}$. First note that the relation only needs to be proven for characters $\chi_{w}$ where $w$ is a word of length $\leq 2$ as the general formula would then follow from Lemma A.3. The relation $\eta\circ S=-\eta$ implies that $\eta+\overline{\eta}\equiv 0$ on the central algebra and therefore, we have
$L(ab)=\varepsilon(a)L(b)+L(a)\varepsilon(b)-\langle\eta(a),\eta(b)\rangle,\quad a,b\in\mathcal{O}(U_{N}^{+})_{0}.$
It follows from a quick computation that $L$ is of the required form for $\alpha:=-L(\chi_{\circ})$ and $\beta:=\|\eta(\chi_{\circ})\|^{2}/2$. ∎
We conclude by explaining why we restrict our study to real-valued (centralized) Gaussian generating functionals on $U_{N}^{+}$. If we were to consider the imaginary part of $\alpha$, convergence may not be guaranteed unless a bound exists that links the real and imaginary parts, similarly to the relation between $\alpha$ and $\beta$. However, we emphasize that, by introducing a drift to the generating functional, the imaginary part can always be disregarded.
@@
$D(ab)=\varepsilon(a)D(b)+D(a)\varepsilon(b),\quad a,b\in\mathcal{O}(U_{N}^{+})_{0}.$
@@
$[D_{H},D_{K}]=D_{H}\star D_{K}-D_{K}\star D_{H}=D_{[H,K]}.$
This shows that drifts correspond to the classical part of Gaussian generating functionals. By adding such a drift, we can always assume a centralized Gaussian generating functional to be real-valued.

### v3

@@ 7.2. Computation of the limit profile
[FFS23, Prop 2.7] Let $L$ be a Gaussian generating functional on $\mathbb{G}=O_{N}^{+}$ or $U_{N}^{+}$. Then, for any $x_{1},\cdots,x_{n}\in\mathcal{O}(\mathbb{G})$ we have
$L(x_{1}\cdots x_{n})=\sum_{1\leq i<j\leq n}L(x_{i}x_{j})\varepsilon(x_{0}\cdots\check{x}_{i}\cdots\check{x}_{j}\cdots x_{n})-(n-2)\sum_{j=1}^{n}L(x_{j})\varepsilon(x_{1}\cdots\check{x}_{j}\cdots x_{n}).$
@@
    $\chi_{w}\longmapsto-\left(\ell(w)\alpha+\big(p(w)-q(w)\big)^{2}\beta+\big(p(w)-q(w)\big)\zeta\right)N^{\ell(w)-1},\quad w\hbox{ word on }\{\lozenge,\blacklozenge\},$
    for some $\alpha,\beta\geq 0$, and $\zeta\in\mathrm{i}\mathbb{R}$.
@@
Let us first note that any Gaussian cocycle $\eta$ defined on a compact quantum group satisfies the relation $\eta\circ S+\eta=0$ (see the beginning of the proof of [FFS23, Thm 3.11]).
Let $(\rho,\eta,L)$ be a Gaussian Schürmann triple on $O_{N}^{+}$. The relation $\eta\circ S=-\eta$ implies that $\eta\equiv 0$ on the central algebra and therefore:
$L(xy)=\varepsilon(x)L(y)+L(x)\varepsilon(y),\quad x,y\in\mathcal{O}(O_{N}^{+})_{0}.$
@@
Let $(\rho,\eta,L)$ be a Gaussian Schürmann triple on $U_{N}^{+}$. The relation $\eta\circ S=-\eta$ implies that $\eta+\overline{\eta}\equiv 0$ on the central algebra and therefore, we have
$L(xy)=\varepsilon(x)L(y)+L(x)\varepsilon(y)-\langle\eta(x),\eta(y)\rangle,\quad x,y\in\mathcal{O}(U_{N}^{+})_{0}.$
It follows from a quick computation that $L$ is of the required form for words of length $\leq 2$ with
$\alpha=-\Re L(\chi_{\lozenge})-\frac{\|\eta(\chi_{\lozenge})\|^{2}}{2N},\quad\beta=\frac{\|\eta(\chi_{\lozenge})\|^{2}}{2N}\zeta=-\mathrm{i}\Im L(\chi_{\lozenge}).$
Let $w=w_{1}\cdots w_{n}$ be a word of length $n\geq 3$. To ease notations, we set $\ell:=\ell(w)$, $p:=p(w)$ and $q:=q(w)$. By Lemma A.3, we compute
$L(w)$ $=\sum_{1\leq i<j\leq n}L(\chi_{w_{i}w_{j}})N^{n-2}-(n-2)\sum_{i=1}^{n}L(\chi_{w_{i}})N^{n-1}$
$=\left(pqL(\chi_{\lozenge\blacklozenge})+\frac{p(p-1)}{2}L(\chi_{\lozenge\lozenge})+\frac{q(q-1)}{2}L(\chi_{\blacklozenge\blacklozenge})\right)N^{n-2}$
$\,\,\,\,\,\,-(n-2)\left(pL(\chi_{\lozenge})+qL(\chi_{\blacklozenge})\right)N^{n-1}$
$=-\Big(n(n-1)-n(n-2)\Big)\alpha N^{n-1}$
$\,\,\,\,\,\,-\Big(2p(p-1)+2q(q-1)-n(n-2)\Big)\beta N^{n-1}$
$\,\,\,\,\,\,-\Big(p(p-1)+q(q-1)-(n-2)(p-q)\Big)\zeta N^{n-1}$
$=-\Big(n\alpha+\big(p-q\big)^{2}\beta+\big(p-q\big)\zeta\Big)N^{n-1}.$
Which proves the result. ∎
We conclude by explaining why we restrict our study to real-valued (centralized) Gaussian generating functionals on $U_{N}^{+}$. If we were to consider the imaginary part, convergence may not be guaranteed unless a bound exists that links the real and imaginary parts, similarly to the relation between $\alpha$ and $\beta$. However, we emphasize that, by introducing a drift to the generating functional, the imaginary part can always be disregarded.
@@
$D(xy)=\varepsilon(x)D(y)+D(x)\varepsilon(y),\quad x,y\in\mathcal{O}(U_{N}^{+})_{0}.$
@@
$D_{H},D_{K}]=D_{H}\star D_{K}-D_{K}\star D_{H}=D_{[H,K]}.$
This shows that drifts correspond to the classical part of Gaussian generating functionals. By adding such a drift111111a centralized drift $D_{H}$ corresponds to a centralized generating functional as described in Proposition A.4 with $\alpha=\beta=0$ and $\zeta=\operatorname{Tr}H$., we can always assume a centralized Gaussian generating functional to be real-valued.
Remark A.6.
Note that every pair $(\alpha,\beta)\in\mathbb{R}_{+}^{2}$ defines a generating functional $L_{\alpha\beta}$ in the sense of Formula (3.2). Indeed, the functionals with $\beta=0$ are obtained by composing a Brownian motion on $O_{N}^{+}$ (see Equation (3.1) with $\nu=0$) with the quotient map $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(O_{N}^{+})$ and the conditional expectation $\mathbb{E}:\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})_{0}$, while the ones with $\alpha=0$ are obtained by following the classification of centralized Gaussian generating functionals on $U_{N}^{+}$ in [FFS24] and considering the case where the construction is made with scalar matrices. All such functionals exist, as the set of all generating functionals forms a convex cone.